# 3D-LFM: Lifting Foundation Model

### v1

@@ 4.2 Benchmark: Object-Specific Models
We benchmark our 3D Lifting Foundation Model (3D-LFM) against leading specialized methods for human body, face, and hands categories. Our model outperforms these specialized methods, showcasing its broad utility without the need for object-specific tailoring, highlighting the versatility of its object-agnostic architecture. For our evaluation, we utilize the H3WB dataset [32], a recent benchmark for diverse whole-body pose estimation tasks. This dataset is especially valuable for its inclusion of multiple object categories and for providing a comparative baseline against methods like Jointformer [14], SimpleBaseline [16], and CanonPose [23]. We followed the H3WB’s recommended 5-fold cross-validation and engaged with the benchmark’s authors to obtain results on the hidden test set. Our reported metrics in Fig. 3 and Table 2 include PA-MPJPE and MPJPE, with the test set performance numbers provided directly by the H3WB team, ensuring that our superior results are verified by an independent third-party.
Performance: Our 3D-LFM demonstrates a notable enhancement in accuracy over baseline methods, as outlined in the referenced figure and table. It excels across all categories, including whole-body, body, face, and hands, as shown in Fig.3 and detailed in Tab.2. With a lean architecture of only $5$ million parameters—significantly fewer than comparative baseline approaches, our model achieves rapid convergence and can be trained efficiently on a single NVIDIA A100 GPU within hours, emphasizing our model’s efficiency and scalability evident by robust performance across varied object categories.
The results affirm 3D-LFM as a versatile and potent foundational model for diverse 2D to 3D lifting tasks, outpacing specialized category-specific models. These findings, showcased in Fig. 3, validate our model’s ability to leverage inter-category learning, potentially paving the way for broader applications across an expanded range of object categories.
@@
Table 2: Performance evaluation of 3D pose estimation models on H3WB and validation datasets showing MPJPE in millimeters. Our method demonstrates leading accuracy across multiple object categories without the need for object-specific designs.
@@ 4.3 Towards foundation model
In this section, we demonstrate the 3D Lifting Foundation Model (3D-LFM) as a versatile foundation for diverse 2D-3D lifting scenarios. Our model is capable of handling various object categories and navigating data imbalance challenges. In this subsection, we explore the 3D-LFM’s scalability and its potential for combined dataset training in Sec. 4.3.1, its proficiency in generalizing to unseen object categories and its adaptability in transferring rig configurations in Sec. 4.3.2. These studies validate the 3D-LFM’s role as a foundation model, capable at leveraging diverse data without requiring specific configurations, thus simplifying the 3D lifting process for varied joint setups.
We show 3D-LFM’s capability of handling 2D-3D lifting for $30+$ object categories within the single model, confirming the model’s capability to manage imbalanced datasets representative of real-world scenarios as shown in Fig. 0 Overview: (a) This figure shows the 3D-LFM’s ability in lifting 2D landmarks into 3D structures across an array of over 30 diverse categories, from human body parts, to a plethora of animals and everyday common objects. The lower portion shows the actual 3D reconstructions by our model, with red lines representing the ground truth and blue lines showing the 3D-LFM’s predictions. (b) This figure displays the model’s training data distribution on a logarithmic scale, highlighting that inspite of 3D-LFM being trained on imbalanced datasets, it preserves the performance across individual categories.. With a comprehensive range of human, hand, face, inanimate objects, and animal datasets, the 3D-LFM is proven to be adaptable, not necessitating category-specific adjustments. The subsequent subsections will dissect these attributes further, discussing the 3D-LFM’s foundational potential in the 3D lifting domain.
@@ 4.3.1 Combined Dataset Training
This study evaluates the 3D-LFM’s performance on isolated datasets against its performance on a comprehensive combined dataset. Initially, the model was trained separately on animal-based supercategory datasets—specifically OpenMonkey[1] and Animals3D[27]. Subsequently, it was trained on a merged dataset encompassing a broad spectrum of object categories, including Human Body-Based datasets like AMASS and Human 3.6 [7], Hands-Based datasets such as PanOptic Hands [9], Face-Based datasets like BP4D+[29], and various Inanimate Objects from the PASCAL3D+ dataset[26], along with the previously mentioned animal datasets. Isolated training resulted in an average MPJPE of $\mathbf{21.22}$ mm, while the integrated training method notably reduced the MPJPE to $\mathbf{12.5}$ mm on the same animal supercategory validation split. This significant improvement confirms the 3D-LFM’s potential as a pretraining framework and underscores its capacity to adapt to and generalize from diverse and extensive data collections.
@@ 4.3.2 OOD generalization and rig-transfer:
We evaluated the 3D-LFM’s generalization ability on object categories and rig configurations unseen during training. The model’s proficiency is demonstrated through its accurate 3D reconstructions of the “Cheetah” category from the Acinoset dataset [10]—distinct from the commonly seen species in the Animal3D dataset [27]. A similar generalization is observed when the model, trained without the “Train” category from PASCAL3D+[26], accurately reconstructs its 3D structure. These findings confirm the model’s robust out-of-distribution (OOD) generalization capability, as qualitatively illustrated in Fig. 4.
Additionally, the 3D-LFM showcased its capability in transferring rig configurations between datasets. For example, training on a $17$-joint Human3.6M dataset [7] and testing on a $15$-joint Panoptic Studio dataset [9] led to accurate 3D reconstructions, despite the disparity in joint numbers and arrangements. These results emphasize the foundational model’s adaptability, critical for processing diverse human data sets with varying rig setups. For a more thorough validation, we direct readers to the ablation section, where extensive qualitative visualizations in Fig. 6 and quantitative reports in Sec. 4.4.3 further substantiate the 3D-LFM’s OOD generalization and rig transfer efficacy.
Figure 4: This figure illustrates 3D-LFM’s proficiency in OOD 2D-3D lifting, effectively handling new, unseen categories from Acinoset [10] PASCAL3D+ [26] with varying joint arrangements, demonstrating its strong generalization capabilities.
@@ 4.4 Ablation
We conduct ablation studies to dissect the 3D-LFM’s design choices and their respecitve contributions. The experiments in Sec. 4.4.1 and Sec. 4.4.2 are carried out on the Human3.6M benchmark [7] and a mixture of Animal3D [27], Human3.6M, and face datasets [9, 29] are used in Sec. 4.4.3 to emphasize the scalability and OOD generalization properties of TPE.
@@ 4.4.1 Procrustean Transformation
The Procrustean approach within 3D-LFM refines learning to the deformable components of objects. Utilizing an OnP solution described in Sec. 3.2, the model focuses on learning deformable shapes in the canonical space and ignoring rigid transformations. The faster convergence and reduced MPJPE error, evident by the gap between blue and orange lines in Fig. 5 (a) validates the benefits of Procrustean transformation, confirming its relevance in the architecture.
Figure 5: (a) Comparing attention strategies in 3D-LFM. The combined local-global approach with OnP alignment surpasses other configurations in MPJPE reduction over 100 epochs on the Human3.6M validation split. (b) rapid convergence and efficiency of the TPE approach compared to the learnable MLP
@@
On the same validation split, our analysis indicates that merging local (GA) and global attention (MHSA) leads to the best performance. Isolating either attention mechanism does not match the performance of their combined use, as shown by the green and red lines in as shown in Fig. 5 (a). The combination of GA with MHSA, especially when aligned with OnP, delivers best results and faster convergence. These findings from our ablation studies validate the architectural decisions behind 3D-LFM.
Our hybrid attention mechanism, which fuses connectivity-aware local graph attention (GA) with the broader scope of global multi-head self-attention (MHSA)—emerges as a more optimal strategy. Sole reliance on either GA or MHSA results in suboptimal performance as evidenced by green and red lines as compared to blue line in Fig. 5 (a). This hybrid attention, paired with Procrustes (OnP) alignment, not only minimizes MPJPE error but also makes model convergence faster, reinforcing the architectural decisions underlying the 3D-LFM.
@@ 4.4.3 Tokenized Positional Encoding:
This ablation study covers the impact of Tokenized Positional Encoding (TPE) in 3D-LFM, which leverages analytical Random Fourier Features for positional information, retaining the permutation equivariance and scalability essential for a foundational model.
Data imbalance study: Our experiments on the Animal3D dataset [27], with a focus on the underrepresented hippo supercategory, reveal TPE’s efficacy in dealing with imbalanced dataset. With a nominal model size of $128$ dimensions, TPE improved MPJPE performance by $\mathbf{3.27}\%$ over the learnable MLP approach, and this performance kept improving as we kept increasing the model size from $128$ to $512$ and finally to $1024$ where the performance improvement kept improving from $\mathbf{3.27}\%$ to $\mathbf{12.28}\%$, and finally to $\mathbf{22.02}\%$ highlighting its capability to handle data imbalance with varying model sizes. More importantly, for all these trials one important thing we noticed is the statistically signifcant faster convergence shown by TPE over learnable MLP approach as evident in Fig. 5 (b).
@@
Rig Transfer Generalization: Rig transfer experiments further solidify TPE’s efficacy, with a $\mathbf{12}\%$ improvement when transferring from a $17$-joint [7] to a $15$-joint rig [9] and an improvement of $\mathbf{23.29}\%$ when we trained on $15$-joint rig and tesed on $17$-joint rig. Training on a $52$-joint facial rig from [9] and testing on an $83$-joint from [29] yielded a $\mathbf{52.3}\%$ improvement over the MLP baseline, validating TPE’s robustness in complex OOD scenarios. Fig. 6 qualitatively shows TPE’s performance, where the our model with TPE is able to generalize to unseen joint configurations. This ablation section touches upon both quantitative and visual studies of TPE and conveys that TPE is a pivotal compoenent for scalable 2D-3D lifting task.
@@ 5 Discussion and Conclusion
The proposed 3D-LFM marks a significant leap in 2D-3D lifting, showcasing scalability and adaptability, addressing data imbalance, and generalizing to new data categories. Its cross-category knowledge transfer and rare shape variation handling requires further investigation to enhance robustness. 3D-LFM’s efficiency is demonstrated by achieving results comparable to leading methods on [32] benchmark as well as its proficiency in out-of-distribution (OOD) scenarios on limited computational resources. For training duration and computational details, please refer to the supplementary materials. This work not only pushes the boundaries of current 2D-3D lifting, but also establishes a scalable framework for future 3D pose estimation and 3D reconstruction models. In summary, the 3D-LFM creates a flexible, universally applicable model for 3D reconstruction from 2D data, paving the way for diverse applications that requires accurate 3D reconstructions from 2D inputs.
Acknowledgement: We extend our gratitude to Ian R. Fasel, Tim Clifford, Javier Movellan, Matthias Hernandez, Matthias Schroeder, and Akshay Subramanya of Apple for their insightful discussions.
@@ I Training Details
The 3D Lifting Foundation Model (3D-LFM), as detailed in Sec. 4.3.1 was trained across more than $30$ diverse categories on a single NVIDIA A100 GPU. This dataset consisted of over 18 million samples, with data heavily imbalanced as shown in Fig. 0 Overview: (a) This figure shows the 3D-LFM’s ability in lifting 2D landmarks into 3D structures across an array of over 30 diverse categories, from human body parts, to a plethora of animals and everyday common objects. The lower portion shows the actual 3D reconstructions by our model, with red lines representing the ground truth and blue lines showing the 3D-LFM’s predictions. (b) This figure displays the model’s training data distribution on a logarithmic scale, highlighting that inspite of 3D-LFM being trained on imbalanced datasets, it preserves the performance across individual categories.. This training setup highlights the model’s practicality, with mixed datasets having imbalance within them. 3D LFM’s transformer block, shown in lower half of Fig. 1 consists of $4$ layers, with hidden dimensions (post TPE dimension) ranging from $64$ to $512$ and head counts between $4$ and $8$. These parameters were adjusted based on the dataset scale in various experiments of Sec. 4. GeLU activations were employed for non-linearity in the feedforward layers. The training process was guided by a ReduceLROnPlateau scheduler with a starting learning rate of 0.001 and a patience of $20$ epochs. An early stopping mechanism was implemented, halting training if no improvement in MPJPE was noted for $30$ epochs, ensuring efficient and optimal performance. This training approach enabled 3D-LFM to surpass leading methods in 3D lifting task proposed by H3WB benchamark [32].
Experimentation in Section 4.2: For the benchmark experiment of H3WB [32], we adopted a 5-fold cross-validation as directed by the benchmark authors. We trained two model variants: one with OnP solution (predicting shapes, $\mathbf{S}_{c}$ in canonical frame) and one without the OnP solution (predicting shapes directly in the reference frame, $\mathbf{S}_{p}$). We aligned the canonical frame predictions from the OnP-based model to those from the non-OnP based model, which ensured that the submitted 3D reconstructions have the rotation and scales adjusted to the predicted reference frame.
II Interactive Visualization with 3D-LFM
Figure 7: Example OOD inferences by 3D-LFM on MBW dataset [5]: The red dots overlaying the images indicate the input 2D keypoints, while the blue stick figures below each image shows the predicted 3D structure viewed from two distinct camera viewing angles. Accompanying video results further illustrate these predictions with a dynamic 360-degree camera animation, providing a detailed view of the model’s performance.
The 3dlfm_visualization.ipynb notebook provided on our project page provides an in-depth interactive exploration of the 3D-LFM’s model predictions – 3D structures from input 2D landmarks. This interactive notebook shows the model’s adaptability across standard validation datasets (shown in Fig. 0 Overview: (a) This figure shows the 3D-LFM’s ability in lifting 2D landmarks into 3D structures across an array of over 30 diverse categories, from human body parts, to a plethora of animals and everyday common objects. The lower portion shows the actual 3D reconstructions by our model, with red lines representing the ground truth and blue lines showing the 3D-LFM’s predictions. (b) This figure displays the model’s training data distribution on a logarithmic scale, highlighting that inspite of 3D-LFM being trained on imbalanced datasets, it preserves the performance across individual categories.) and its adaptability to challenging Out-Of-Distribution (OOD) scenarios, particularly from in-the-wild captures provided by the Acinoset datasets [10] and MBW datasets [5].
II.1 Key Features
Category Selection: Enables selection from diverse object categories, including standard and OOD. Data Management: Automates the loading and visualization of 2D inputs, 3D ground truth, and predictions. Frame Exploration: Facilitates detailed visualization of specific frames with the option to overlay 2D input data on images and viewing its corresponding 3D predictions simultaneously.
II.2 Usage and examples
To use the notebook, simply select an object category and frame number (automated within). It will display the model’s 3D and 2D predictions, alongside available ground truth data. For OOD categories, there’s an option to include pseudo ground truth labels, offering a deeper insight into the model’s performance in diverse scenarios.
As shown in Fig. 7, the 3D-LFM handles complex OOD data, generating predictions in-the-wild with reasonable accuracy. Unlike traditional approaches requiring extensive bootstrapping to generate labels for tail-end distribution categories, our 3D-LFM generates 3D predictions in a single inference step, emphasizing its foundation model capabilities. These include in-the-wild reconstructions of various wildlife captured in-the-wild, as showcased in Fig. 7. This figure, alongside additional visualizations within the notebook shows that 3D-LFM is capable of handling multiple categories within a single unified model. It particularly handles OOD scenarios reasonably well – providing accurate 2D-3D lifting performance on in-the-wild captures from datasets like Acinoset [10] and MBW [5]. These findings confirm the 3D-LFM’s ability to utilize its foundational attributes for accurate 2D-3D lifting, even across previously unseen categories.
III Limitations
Category Misinterpretation: The 3D-LFM exhibits remarkable performance generalizing across object categories. However, it can encounter difficulties when extreme perspective distortions cause 2D inputs to mimic the appearance of different categories. For example, in the case of a tiger viewed from an atypical angle, the model’s output may resemble the structure of a primate due to similar 2D keypoint configurations induced by perspective effects, as shown in Fig. 8 (c). This confusion stems from the model’s reliance on the geometric arrangement of keypoints, which can be deceptive under extreme perspectives.
Depth Ambiguity: Another challenge arises from the inherent depth ambiguity in single-frame 2D to 3D lifting. The model can misinterpret the spatial arrangement of limbs, as evident by Fig. 8 (a), where a monkey’s leg appears to extend backward rather than forward. Similarly, Fig. 8 (b) displays an instance where the model perceives the monkey’s head orientation incorrectly. Such depth ambiguities highlight the need for more sophisticated depth cues in single-view 3D reconstruction.
We propose integrating appearance cues, such as those provided by DINOv2 features, to enhance depth perception and category differentiation. This would provide additional context to the geometric data, aiding in resolving ambiguities, as demonstrated by the OOD examples in Fig. 8, derived from challenging in-the-wild captures [5]. This integration aims to improve the fidelity of 3D reconstructions in scenarios with complex visual information.

### v2

@@ 4.2 Benchmark: Object-Specific Models
Next, we benchmark 3D-LFM against leading specialized methods for human body, face, and hands categories. Our model outperforms these specialized methods, showing multi-category learning without the need for category (landmark) semantics. For this study, we evaluate on H3WB dataset [32], a recent benchmark for diverse whole-body pose estimation tasks. This dataset is valuable for its inclusion of multiple object categories and for providing a comparative baseline against methods such as Jointformer [14], SimpleBaseline [16], and CanonPose [23]. Following H3WB’s recommended 5-fold cross-validation and submitting the evaluations to benchmark’s authors, we report results on the hidden test set. The results shown in Fig. 3 and Table 1 include PA-MPJPE and MPJPE, with test set performance numbers provided directly by the H3WB team, ensuring that our results are verified by an independent third-party.
@@
Table 1: Quantitative performance on H3WB: Our method demonstrates leading performance across multiple object categories without the need for object-specific designs.
@@ 4.3 Towards foundation model
In this section, we highlight 3D-LFM’s role as a foundational model for varied 2D-3D lifting, capable in managing multiple object types and data imbalances. In this subsection, we explore 3D-LFM’s scalability for collective dataset training (Sec.4.3.1), its generalization to new categories and rig transfer capabilities (Sec.4.3.2). These studies validate the 3D-LFM’s role as a foundation model, capable at leveraging diverse data without requiring specific configurations, thus simplifying the 3D lifting process for varied joint setups.
We start this investigation by showing the capability 3D-LFM in handling 2D-3D lifting for $30+$ object categories within the single model, confirming the model’s capability to manage imbalanced datasets representative of real-world scenarios as shown in Fig. 0 Overview: (a) This figure shows the 3D-LFM’s ability in lifting 2D landmarks into 3D structures across an array of over 30 diverse categories, from human body parts, to a plethora of animals and everyday common objects. The lower portion shows the actual 3D reconstructions by our model, with red lines representing the ground truth and blue lines showing the 3D-LFM’s predictions. (b) This figure displays the model’s training data distribution on a logarithmic scale, highlighting that inspite of 3D-LFM being trained on imbalanced datasets, it preserves the performance across individual categories.. With a comprehensive range of human, hand, face, inanimate objects, and animal datasets, the 3D-LFM is proven to be scalable, without requiring category-specific adjustments. The subsequent subsections will dissect these attributes further, discussing the 3D-LFM’s foundational potential in the 3D lifting domain.
@@ 4.3.1 Combined Dataset Training
This study evaluates the 3D-LFM’s performance on isolated datasets against its performance on a combined dataset. Initially, the model was trained separately on animal-based supercategory datasets: specifically OpenMonkey[1] and Animals3D[27]. Subsequently, it was trained on a merged dataset containing a broad spectrum of object categories, including Human Body-Based datasets such as AMASS [15] and Human 3.6 [7], Hand-Based datasets such as PanOptic Hands [9], Face-Based datasets like BP4D+[29], and various Inanimate Objects from the PASCAL3D+ dataset[26], along with previously mentioned animal datasets. Isolated training resulted in an average MPJPE of $\mathbf{21.22}$ mm, while the combined training method significantly reduced MPJPE to $\mathbf{12.5}$ mm on the same animal supercategory validation split. This improvement confirms the potential of 3D-LFM as a pre-training framework and underscores its ability to adapt and generalize from diverse and extensive data collections.
Dataset Selection Rationale: We selected animal-based supercategory datasets to demonstrate combined training’s impact on underrepresented categories. We observed greater performance improvements in smaller, unbalanced datasets (as exemplified by PASCAL3D+: from $\mathbf{4.31}$ mm to $\mathbf{1.1}$ mm and OpenMonkey: from $\mathbf{19.45}$ mm to $\mathbf{9.59}$ mm) compared to larger datasets with sufficient balance among categories. Consequently, we see minimal gains in more balanced, larger datasets like AMASS (from $\mathbf{1.67}$ mm to $\mathbf{1.66}$ mm), underscoring the utility of combined training for enhancing performance in underrepresented and long-tail categories.
@@ 4.3.2 OOD generalization and rig-transfer:
We evaluate 3D-LFM’s generalization to unseen object categories and rig configurations. Its accuracy is highlighted by successful 2D-3D lifting reconstructions of the “Cheetah” from Acinoset [10], which is not included in the typical Animal3D dataset [27], and the “Train” category from PASCAL3D+[26], absent during training. Qualitative reconstructions are shown in Fig. 4, along with the quantitative results in Tab.2 for above categories as well as in-the-wild category like a Chimpanzee from the MBW dataset [5] – which illustrates model’s strong OOD generalization and capability to handle in-the-wild data.
Table 2: Quantitative evaluation for OOD scenarios
Category OOD (mm) In-Dist. (mm)
Cheetah [10] $26.59$ $10.16$
Train [26] $6.88$ $5.71$
Chimpanzee [5] $52.05$ $42.65$
Additionally, we show 3D-LFM’s capability in transferring rig configurations between datasets, embodying the concept of generic geometry learning. By training on a $17$-joint Human3.6M dataset [7] and testing on a $15$-joint Panoptic Studio setup [9], our model gives accurate 3D reconstructions despite variations in joint arrangements. This capability is particularly interesting for its efficiency in utilizing data from multiple rigs of the same object, and underscores the model’s adaptability, a cornerstone in processing diverse human datasets. It aligns with the broader community’s interest in versatile geometry learning, which makes these findings especially compelling. For a more thorough validation, we direct the reader to the ablation section, where qualitative visuals (Fig. 6) and quantitative analysis (Sec. 4.4.3) further highlight 3D-LFM’s OOD generalization and rig transfer efficacy.
Figure 4: Generalization to unseen data: Figure showing 3D-LFM’s proficiency in OOD 2D-3D lifting, effectively handling new, unseen categories, and rig generalization from Acinoset [10] PASCAL3D+ [26], and Panoptic studio [9] with varying joint arrangements in top row. The bottom row presents in-the-wild data from the MBW dataset [5], with red dots indicating input keypoints and blue stick figures showing the model’s 3D predictions from different angles.
@@ 4.4 Ablation
In our ablation studies, we evaluate the 3D-LFM’s design elements and their individual contributions to its performance. Detailed experiments on the Human3.6M benchmark [7] and a blend of other datasets including Animal3D [27] and facial datasets [9, 29] were carried out to ablate the role of Procrustean transformation, hybrid attention mechanisms, and tokenized positional encoding (TPE) in enabling the model’s scalability and out-of-distribution (OOD) generalization.
@@ 4.4.1 Procrustean Transformation
3D-LFM’s fusion of the procrustean approach, a first in transformer-based lifting frameworks, concentrates on deformable object components, as outlined in Sec.3.2. By focusing on shape within a standard canonical reference frame and avoiding rigid body transformations, we see faster learning and a decreased MPJPE, as evident by the gap between blue and orange lines in Fig. 5 (a) suggests. This fusion is crucial for learning 3D deformations, while utilizing transformers’ equivariance. These findings suggest that even for transformers, avoiding rigid transformations’ learning aids convergence, most notably with imbalanced datasets.
Figure 5: (a) Comparing attention strategies in 3D-LFM. The combined local-global approach with procrustean alignment surpasses other configurations in MPJPE reduction over 100 epochs on the Human3.6M validation split. (b) rapid convergence and efficiency of the TPE approach compared to the learnable MLP
@@
In evaluating 3D-LFM’s attention strategies, our analysis on the same validation split as above demonstrates the superiority of a hybrid approach combining local (GA) and global (MHSA) attention mechanisms. This integration, particularly when complemented by Procrustean alignment, significantly enhances performance and accelerates convergence, as evidenced in Fig. 5 (a). The distinct advantage of this hybrid system validates our architectural choices, showcasing its efficiency in reducing MPJPE errors and refining model training dynamics.
@@ 4.4.3 Tokenized Positional Encoding:
This ablation study assesses the impact of Tokenized Positional Encoding (TPE), which uses analytical Random Fourier Features for encoding positional information. This study examines TPE’s influence on model performance in scenarios of data imbalance and rig transfer generalization.
Table 3: Impact of TPE on Data Imbalance and Rig Transfer
Study Experiment Model Size Improvement (%) Data Imbalance Underrepr. category (Hippo)  [27] $128$ $3.27$ $512$ $12.28$ $1024$ $22.02$ Rig Transfer $17$ [7]- to $15$ [9]-joint N/A $12$ $15$ [9]- to $17$ [7]-joint $23.29$ $52$ [9]- to $83$ [29]-joint $52.3$
Data imbalance study: When tested on the underrepresented hippo category from the Animal3D dataset [27], TPE based model showed a $3.27$% improvement in MPJPE over the baseline MLP with a $128$-dimensional model performance as evident in first row of Tab. 3. This improvement grew with the model size. These results highlight TPE’s scalability and its faster convergence, especially relevant in imbalanced, OOD scenarios as detailed in Fig. 5 (b). The observed performance boosts suggest that TPE’s analytical nature might be more suited to adapting to novel data distributions. Increasing model size amplifies TPE’s benefits, hinting that its fixed analytical approach more adeptly handles OOD intricacies compared to learnable methods like MLPs, which may falter in such situations.
@@
Rig transfer study: Our rig transfer analysis, summarized in Table 3, showcases TPE’s adaptability and effectiveness over the MLP baseline across different joint configurations and rig scenarios, with improvements up to $52.3$%. These findings, particularly the significant performance boost in complex rig transfers, underscore TPE’s robustness in OOD contexts. Figure 6 visually highlights the qualitative differences between TPE and MLP approaches in a rig transfer scenario, where the model trained on a $17$-joint [7] configuration is tested on a $15$ joint [9] setup.
@@ 5 Discussion and Conclusion
The proposed 3D-LFM marks a significant leap in 2D-3D lifting, showcasing scalability and adaptability, addressing data imbalance, and generalizing to new data categories. Its cross-category knowledge transfer requires further investigation and handling of inputs with different perspectives could act as potential limitations. 3D-LFM’s efficiency is demonstrated by achieving results comparable to leading methods on [32] benchmark as well as its proficiency in out-of-distribution (OOD) scenarios on limited computational resources. For training duration and computational details, please refer to the supplementary materials. This work establishes a baseline framework for future 3D pose estimation and 3D reconstruction models. In summary, the 3D-LFM creates a universally applicable model for 3D reconstruction from 2D data, paving the way for diverse applications that requires accurate 3D reconstructions from 2D inputs.
Acknowledgement: We extend our gratitude to Ian R. Fasel, Tim Clifford, Javier Movellan, and Matthias Hernandez of Apple for their insightful discussions.
@@ I Training Details
Model parameters: In the architecture of 3D-LFM, the transformer block consists of four layers, with the hidden dimensions and head counts tailored to the dataset scale. Specifically, for datasets exceeding $10,000$ frames, we used a model dimension of $512$ and a head count of $8$. Datasets with frame counts ranging from $1,000$ to $10,000$ were assigned model dimensions of $256$ with $4$ heads. For smaller datasets, consisting $1$ to $1,000$ frames, we employed a more compact model with dimensions set at $64$ and head counts maintained at $4$. This gradation in model complexity ensures a balanced approach, aligning the model capacity with the dataset size, which is particularly critical for achieving computational efficiency and avoiding overfitting on smaller datasets.
Optimizer and scheduler: GeLU activations were employed for non-linearity in the feedforward layers. The training process was guided by a ReduceLROnPlateau scheduler with a starting learning rate of 0.001 and a patience of $20$ epochs. An early stopping mechanism was implemented, halting training if no improvement in MPJPE was noted for $30$ epochs, ensuring efficient and optimal performance. This training approach enabled 3D-LFM to surpass leading methods in 3D lifting task proposed by H3WB benchamark [32].
II Limitations
Perspective-Induced Misinterpretations: The 3D-LFM demonstrates a significant capability in generalizing across object categories. However, it can encounter difficulties when extreme perspective distortions cause 2D inputs to mimic the appearance of different categories. For example, unusual viewing angles can cause a tiger to be misconstrued as a primate, illustrated in Fig. 7 (c), due to the similarity in 2D keypoint configurations. Similarly, depth ambiguities can result in incorrect interpretations of spatial arrangements, such as a monkey’s leg extending backward instead of forward (Fig. 7 (a)) or misperceiving the orientation of a monkey’s head (Fig.  7 (b)). These limitations highlight the complexities of single-frame 2D to 3D lifting, where the model’s reliance on geometric keypoint arrangements can be deceptive under certain perspectives. Enhanced depth cues and perspective-aware mechanisms are necessary for more accurate single-view 3D reconstructions, pointing towards future directions to integrate additional appearance or visual features based contextual information for resolving these ambiguities.